OpenStudy (anonymous):

solutions to 6x^2=18x

OpenStudy (e.mccormick):

First, you need to set it to zero. Then find the factors.

OpenStudy (anonymous):

Hope it helps

OpenStudy (anonymous):

@50_cent copy and paste is not correct! You copied and pasted from the website: http://www.geteasysolution.com/6x%5E2-18x=0

OpenStudy (anonymous):

@50_cent thats not even the same equation :(

OpenStudy (anonymous):

@e.mccormick i don't understand. How??

OpenStudy (e.mccormick):

If you look at what I said, it is the set it to 0 part.

OpenStudy (e.mccormick):

\(6x^2=18x\implies 6x^2-18x=18x-18x\implies 6x^2-18x=0\)

OpenStudy (mertsj):

1. Subtract 18x from both sides 2. Factor out the common factor 6x 3. Set each factor equal to 0 4. Solve each equation 5. Write your answers.

OpenStudy (anonymous):



OpenStudy (e.mccormick):

Hehe. We are all saying the same thing. OK.. so, now, do you get the first step?

OpenStudy (anonymous):

lols ikr. Is there more steps? :0 my teacher told me the answer was x=0 x=3 and i just want to find out how she got that xD..and No i dont get it lol

OpenStudy (freethinker):

Solve for x over the real numbers: 6 x^2 = 18 x Move everything to the left hand side. Subtract 18 x from both sides: 6 x^2-18 x = 0 Factor the left hand side. Factor x and constant terms from the left hand side: 6 (x-3) x = 0 Divide both sides by a constant to simplify the equation. Divide both sides by 6: (x-3) x = 0 Solve each term in the product separately. Split into two equations: x-3 = 0 or x = 0 Look at the first equation: Solve for x. Add 3 to both sides: Answer: | | x = 3 or x = 0

OpenStudy (anonymous):

@freethinker ahh :D thank you!!

OpenStudy (anonymous):

\[6 x^2-18 x = 0\] \[x = \frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a }\] \[x = \frac{ -(-18) \pm \sqrt{(-18)^2 - 4(6)(0)} }{ 2(6) }\] \[x = \frac{ 18 \pm \sqrt{324 - 0} }{ 12 }\] \[x = \frac{ 18 \pm \sqrt{324} }{ 12 }\] \[x = \frac{ 18 \pm 18 }{ 12 }\] \[x = \frac{ 18 + 18 }{ 12 }\] \[x = \frac{ 36}{ 12 }\] \[x = 3\] \[x = \frac{ 18 - 18 }{ 12 }\] \[x = \frac{ 0 }{ 12 }\] \[x = 0\]

OpenStudy (freethinker):

that too to complete the square

OpenStudy (e.mccormick):

And there it is with the quadratic formula. Wheee. Yah. What freethinker did is where it all goes from where I was starting. The point here is learning about factors. Know how you sometimes factor a number? Like \(14=7\cdot 2\) or \(125=5\cdot 5\cdot 5\). These principals carry over into algebra. The factors of \(6 x^2-18 x \) are the things that would multiply together to make it. Factoring is un-multiplying, which is not the same as dividing.

OpenStudy (anonymous):

Algerbra is hard .-. ...But that made me understand more @e.mccormick Thanks! :3

OpenStudy (anonymous):

lol im the only one without a medal :( haha

OpenStudy (e.mccormick):

Mertsj didn't get one.... but Mertsj has more than enough medals... =P

OpenStudy (anonymous):

lol yeah he does ;)

OpenStudy (freethinker):

Solve for x over the real numbers: 6 x^2 = 18 x Move everything to the left hand side. Subtract 18 x from both sides: 6 x^2-18 x = 0 Write the quadratic equation in standard form. Divide both sides by 6: x^2-3 x = 0 Take one half of the coefficient of x and square it, then add it to both sides. Add 9/4 to both sides: x^2-3 x+9/4 = 9/4 Factor the left hand side. Write the left hand side as a square: (x-3/2)^2 = 9/4 Eliminate the exponent on the left hand side. Take the square root of both sides: x-3/2 = 3/2 or x-3/2 = -3/2 Look at the first equation: Solve for x. Add 3/2 to both sides: x = 3 or x-3/2 = -3/2 Look at the second equation: Solve for x. Add 3/2 to both sides: Answer: | | x = 3 or x = 0

OpenStudy (anonymous):

woah woah woah dude. you cant just change how you solve it @freethinker now im confused as hell xD

OpenStudy (freethinker):

nawww it's another way of solving it called completing the square.

OpenStudy (anonymous):

6 (x-3) x = 0 Divide both sides by a constant to simplify the equation -------- in here when u divide by 6..what else is getting divided by 6?

OpenStudy (anonymous):

oooh havnt heard of it xD

OpenStudy (freethinker):

both sides of the eQUATION.

OpenStudy (anonymous):

.-. oh

OpenStudy (freethinker):

6 (x-3) x = 0 -------- -- 6 6 6 ON the left-hand side (LHS) cancels out leaving you with (x-3) 0 divided by 6 is 0

OpenStudy (e.mccormick):

Lesson #2 about math... there is more than one way to skin a factoring problem. You just saw three valid solutions. Here is a fourth, it is a graphical one. Where do the lines meet? What are the x values at those points?

OpenStudy (anonymous):

hmm. Interesting.

OpenStudy (e.mccormick):

Find a method that works for you. Get used to it. But remember that you may be required to use other methods in the future. By keeping up on old skills and developing new ones, you can tackle pretty much any problem as the come along. Also, many degrees these days require you take calculus. The vast majority of a calculus class is not calculus, it is trig and algebra and geometry and even arithmetic. So if you are better at these basic skills, you will do better in calculus.

OpenStudy (anonymous):

The best method was the first you gave me . :) ill keep this in mind! i appreciate it thanks.

OpenStudy (e.mccormick):

Exactly, for something this simple, the easy factor is there. For something more complex, there is the big gun of the quadratic formula. If it is for a circle or hyperbola solution, you MUST use completing the square. All are tools, all have a job, use the right tool and it is easy... wrong tool and it can be a lot more work to get the same answer!

OpenStudy (anonymous):

true true. i checked on google and it gave me something totally different!

OpenStudy (e.mccormick):

What did you find there?

OpenStudy (anonymous):

the same answer 50 cent gave me.

OpenStudy (anonymous):

@frann what math course are you taking? Also what math knowledge do you have? For example, general math, pre-algebra, algebra 1, algebra 2, ... etc.

OpenStudy (anonymous):

That is important cuz with that we know till what to show you. Like for example, you might not heard of completing the square since you probably have not gotten there in math.

OpenStudy (anonymous):

Algerbra 1 . But struggling ;-;

OpenStudy (e.mccormick):

There are many good math references out there. I have used many, but when it comes to algebra I do have a favorite. http://www.purplemath.com/modules/index.htm If you match up the name of what you are doing in your class to the topics there, you can usually find a good description. Many are not too long. However, there are complex topics where they take several pages to go over and do multiple examples. I find they have more information than most textbooks I have seen coupled with a practical way of presentation. That way you can learn the techniques, not the answers. If you know the techniques, the answers are there for the picking!

OpenStudy (anonymous):

hmm okaii.

OpenStudy (anonymous):

:) thank yew haha. Some websites out there never really help including my actual math teacher

OpenStudy (e.mccormick):

yah, I looked at that http://www.geteasysolution.com/6x%5E2-18x=0 thing and it is technically right.... but it is presented poorly. You have to know what it means to know how it got there. If you knew what it meant, you would not need the information. So it is worthless!

OpenStudy (anonymous):

@e.mccormick i never said it was wrong, i didn't even take a look at it; in depth. I just said that the user copied and pasted the answer from there lol :P

OpenStudy (anonymous):

The best one ive used was yourteacher.com but after the free trial you had to pay. so i was bummed out :/ It was the ONLY one that i understood

OpenStudy (e.mccormick):

@some_someone Hehe. I am not saying you said it was wrong. I know you meant it was wrong of him to just do that. I am saying that without an explanation, that web site is worthless, which is pretty much what you were pointing out too.

OpenStudy (anonymous):

lol yep :)

OpenStudy (anonymous):

lol how random @Luis_Rivera

OpenStudy (e.mccormick):

@frann One other problem with math teachers is that many forget to tell you that you are in a language class. Mathematics is a language. It has all these terms that mean very specific things in math. If you just keep teaching math like it is counting things on one hand, with no explanation that you are now entering a new language, how will you ever expect these odd bits of grammar and "strange" terms? Let me give you an example. This is actually related to algebra 1, but I bet you would not know a lot of what it says. It is given that \(\{a, b, x, x_1, x_2\}\in \mathbb{R}\) and \(x_1\ne x_2\). If \(x=\{x_1,x_2\}\) for \(ax=b\) then: \(ax_1=b;\, ax_2=b \) \[b=b\\ \implies ax_1= ax_2 \\ \implies x_1= x_2 \forall x_1, x_2\]\(\because x_1\ne x_2,\; ax=b\) must only have one solution. Q.E.D.

OpenStudy (anonymous):

*o* ooh.

OpenStudy (anonymous):

hey @Luis_Rivera :D

OpenStudy (e.mccormick):

how will you ever expect these odd bits of grammar and "strange" terms? ... oops... to make sense was supposed to be in there somewhere.

OpenStudy (anonymous):

its ok ! :D i understood it

OpenStudy (e.mccormick):

What that bottom part means is that \(3x=9\) has only one answer, \(x=3\). But it is what they call a proof. They use things like \(\in\) which is "in" or "is a member of" and \(\implies\) implies.

OpenStudy (anonymous):

its like using symbols :o and meaning and stuff

OpenStudy (e.mccormick):

Like ;P and ;-) and so on. Exactly. It is a language. That language has rules, like what "solutions to \(6x^2=18x\)" means and what it implies about how to get there. If you were sad, you would not send out ;P, you would send ;-; or :( out. What the message means implies things. What the formula or question given menas is the same thing.

OpenStudy (e.mccormick):

menas.... doh. Ignore the lysdexia. Means.

OpenStudy (anonymous):

hahaha menas. well i suppose i can use math as a lanuguage now .But i wont be so good on it.

OpenStudy (e.mccormick):

Well, lets come full circle. I am just talking about a concept so you can help understand why math is almost alien at times. That equation, \(6x^2=18x\) can be read as "where does the curve six ex squared equal or intersect with the line eighteen ex?"

OpenStudy (e.mccormick):

The answer to it stated that way is the picture I linked where they cross. It is just one reading of the language. But understanding the intent leads to all the possible methods of getting that solution.

OpenStudy (anonymous):

ooooh thats how you got the points. Could be represented as (0,3)

OpenStudy (e.mccormick):

Yes, and when you know a language, you know how you can say the same things diferent ways. Hello \(\implies\) Hi \(\implies\) What's up! \(\implies\) Yo \(6x^2=18x\implies 6x^2-18x=0\)

OpenStudy (anonymous):

ooh i see. i have one more problem, if you can help me answer it?

OpenStudy (e.mccormick):

Sure. I am just talking conseptual stuff now because you seemed to need to connect with the math a bit more.

OpenStudy (anonymous):

yess hehe :) . A company builds computers. It costs 6,700 to build 10 computers and 12,00 to build 20 computers. which equation models the cost c(x) as a linear function of the number of computers build?

OpenStudy (e.mccormick):

12,00 or 12,000? The language is the one used by Mathematicians... a strange breed the came out of Philosophers. We are not sure how... probably genetic drift or survival instinct. =P But once Mathematicians started breeding, they infested every country and became deeply vested in all the sciences. So knowing their language is a necessary part of dealing with a large chunk of our world. If you want a truly bizarre language, learn Philosophy itself. There they have had questions documented as having been asked 3000 years ago, and they are still trying to find an answer!

OpenStudy (anonymous):

1200. :) and philosophy itself is confusing enough.

OpenStudy (anonymous):

12,200...OOPS!

OpenStudy (e.mccormick):

ah. K. Now, what are the important numbers and words in that problem. Tell me which ones you see and I'll tell you which ones I see.

OpenStudy (anonymous):

I see 6,700 and 12,200

OpenStudy (e.mccormick):

I see: costs 6,700 to build 10 costs 12,200 to build 20 cost c(x) as a linear function of the number of computers build

OpenStudy (e.mccormick):

What do you think "a linear function" means?

OpenStudy (anonymous):

When the slope is equal? I think :E

OpenStudy (e.mccormick):

Not bad. Sort of. What do you think "a \(\huge \mathrm{line}\)ar function" means? (big hint)

OpenStudy (anonymous):

when the line of the slope is straight?

OpenStudy (e.mccormick):

Yes, some of math is this basic and uncreative. A curve has a slope that changes, and many other things. A line, well, is a line.... it is straight and goes forever, so we deal with little segments we can see and understand. In this case, they want a function that makes a line.

OpenStudy (e.mccormick):

So, if we are looking for the equation of a line, and have these things: costs 6,700 to build 10 costs 12,200 to build 20 And we are told that function is to find the cost c(x), what does that make those numbers?

OpenStudy (anonymous):

it makes it linear?

OpenStudy (e.mccormick):

Line = Linear But I am taling more like: 

OpenStudy (e.mccormick):

So which is the x and which is the y? And do you know how to find the slope of a line from two points?

OpenStudy (anonymous):

yes : \[\frac{ y _{2}-y _{1} }{ x _{2} -x _{1}}\]

OpenStudy (e.mccormick):

OK. So, what is the \((x_1,y_1)\) and what is the \((x_2,y_2)\)?

OpenStudy (e.mccormick):

The Ordinate is ordinary Y or Results. The Abscissa is the x. But those terms confuse people, so we will stick with Y and X axis. We have c(x) on the y axis and x on the... wait for it... x axis. (not surprising it is there.) That makes the x,y points into what?

OpenStudy (anonymous):

(6,700 , 12,200) (10 , 20) ?

OpenStudy (e.mccormick):

That is two outputs together and teo inputs together... not quite right.